# Shape Formation and Locomotion with Joint Movements in the Amoebot Model

### v2

@@ 2.3 Problem Statement and Our Contribution
In this paper, we formalize and extend the joint movement extension proposed by Feldmann et al. [23]. In the following, we provide a proof of concept. For that, we focus on two fundamental problems of MRSs: shape formation and locomotion. We study these problems from a centralized view to explore the limits of the extension.
In the shape formation problem, an MRS has to reconfigure its structure into a given shape. For that, we define meta-modules of rhombical and hexagonal shape. We show that these meta-modules are able to perform various movement primitives of other MRSs, e.g., crystalline atoms, and rectangular/hexagonal metamorphic robots. This allows us to simulate the shape formation algorithms of those models.
@@ 2.4 Related Work
Various MRSs have been defined for different lattices, e.g., [12, 60] utilize the triangular grid, and [3, 22, 49] utilize the Cartesian grid. We will build meta-modules for the Cartesian and hexagonal grid. Note that some MRSs were also physically realized. Examples for MRSs using the triangular grid are hexagonal metamorphic robots [12], HexBots [50], fractal machines [42], and catoms [34]. Examples for MRSs using the Cartesian grid are CHOBIE II [54], EM-Cubes [7], M-Blocks [47], pneumetic celluar robots [28], and XBots [59].
@@
The most common movement primitives for MRSs on the Cartesian grid are the rotation and slide primitives (see Figures 5(b) and 5(c)), and for MRSs on the triangular grid the rotation primitive (see Figure 5(a)). Some models may constrain these movement primitives (see Figure 7). We refer to [2, 27] for a deeper discussion about possible constraints. One way to bypass such constraints is to construct meta-modules, e.g., see [19, 27, 43]. Our meta-modules implement all aforementioned movement primitives without any constraints, i.e., they perform them in place.
@@
Figure 7: Constrained movement primitives. The green modules are moving, respectively. The gray cells must be empty. Both primitives have the same result while the pivoting model requires more free space than the sliding model.
From all aforementioned models, crystalline atoms [48], telecubes [53, 57], and prismatic cubes [58] are the closest to our joint movement extension. In these MRSs, each robot has the shape of a unit square and is able to expand an arm from each side by half a unit. Adjacent robots are able to attach and detach their arms. Similar to our joint movement extension, a robot can move attached robots by expanding or contracting its arms. In contrast to amoebots, a line of robots cannot reconfigure to any other shape since each pair of robots can only have a single point of contact [5]. In order to allow arbitrary shape formations, the robots are combined into meta-modules of square shape that are capable of various movement primitives, e.g., the rotation, slide, contraction, and tunnel primitives (see Figures 5(b), 5(d), 5(c) and 5(e)). Our rhombical meta-modules can implement all these movement primitives. The rhombical shape has two advantages compared to the square shape. First, we can implement further movement primitives that provide us with a simple way to construct a walking structure. Second, we can utilize rhombical meta-modules as a basis for hexagonal meta-modules.
@@ 3 Meta-Modules
In this section, we will combine multiple amoebots to meta-modules. In other models for programmable matter and modular robots, meta-modules have proven to be very useful. For example, they allow us to bypass restrictions on the reconfigurability [19, 56] and to simulate (shape formation) algorithms for other models [4, 46].
In the first two subsections, we will present meta-modules of rhombical and hexagonal shape, respectively. In the third and fourth subsection, we discuss how to increase the stability of the structure during the movements, and how to construct rhombical and hexagonal meta-modules from an arbitrary amoebot structure.
@@ 3.1 Rhombical Meta-Modules
Let $\ell$ be a positive even integer. Our rhombical meta-module consists of $\ell^{2}/2$ uniformly oriented expanded amoebots that we arrange into a rhombus of side length $\ell-1$ (see Figure 8). We obtain a parallelogram of side lengths $\ell-1$ and $\ell/2-1$ if we contract all amoebots (see Figure 7(a)). Note that we have to remove some bonds to perform the contraction. We can expand the parallelogram again by reversing the contractions.
@@
(a) Contraction primitive.
(b) Reorientation primitive.
(c) Realignment primitive.
(d) Rotation primitive.
(e) Slide primitive.
(f) $k$-tunnel primitive.
Figure 8: Movement primitives for rhombical meta-modules. Red meta-modules perform a pull operation, and blue meta-modules a push operation.
There are exactly two possibilities to arrange the uniformly oriented expanded amoebots in a rhombus. By reorienting the amoebots in pairs with the help of handovers, we can reorientate all amoebots within a rhombus (see Figure 7(b)).
@@
Our implementation of the reorientation primitive requires $3$ rounds.
Furthermore, there are three possibilities to align the sides of a rhombus to the axes of the triangular grid. By sliding each second row along its axis to the other side, we can realign the other axis a rhombus is aligned to (see Figure 7(c)). Note that in combination with the reorientation of the amoebots within a rhombus, we are able to align a rhombus with any two axes of the triangular grid.
@@
We can arrange the meta-modules on a rhombical tesselation of the plane if they are all aligned to the same axes (see Figure 0(b)). Note that due to the triangular grid, the meta-modules are not connected diagonally everywhere. Hence, we will only consider meta-modules connected if their sides are connected. In the following, we introduce two movement primitives: the slide and $k$-tunnel primitive. Our implementations of these primitives are similar to the ones for crystalline robots (e.g., [5]) and teletubes (e.g., [57]).
In the slide primitive, we move a meta-module $R_{1}$ along two adjacent substrate meta-modules $R_{2}$ and $R_{3}$ (see Figure 5(c)). We realize the primitive as follows (see Figure 7(e)). We assume that all amoebots are orientated into the movement direction. Otherwise, we apply the reorientation primitive. With respect to Figure 7(e), let $L_{1}$ denote the uppermost layer of $R_{2}$ and $R_{3}$, and $L_{2}$ the second uppermost layer of $R_{2}$ and $R_{3}$. Our slide primitive consists of two rounds. In the first round, we contract all amoebots in $L_{1}$ after removing all bonds between $L_{1}$ and $R_{1}$ except the last one in the movement direction, and all bonds between $L_{1}$ and $L_{2}$ except the last one in the opposite direction. This moves $R_{1}$ into its target position. In the second round, we restore $R_{2}$ and $R_{3}$. For that, we expand $L_{1}$ again after removing all bonds between $L_{1}$ and $R_{1}$, and between $L_{1}$ and $L_{2}$ except the last ones in the movement direction, respectively. This ensures that $R_{1}$ stays in place.
@@
In the $k$-tunnel primitive, we move a meta-module $R$ through a simple path of meta-modules with $k$ corners to the other end (see Figure 5(e)). However, we do not move $R$ directly through the path. Instead, we make use of the following two basic operations. First, by contracting two adjacent meta-modules into parallelograms, we pull one meta-module into the other (see the first round in Figure 7(f)). Second, by expanding two adjacent contracted meta-modules, we push one meta-module out again (see the fourth round in Figure 7(f)).
Now, consider a line of at least 4 meta-modules with two contracted meta-modules at one end. By expanding those two meta-modules and contracting the two meta-modules at the other end, we transfer a meta-module from one end to the other end without changing the length of the line (see the third round in Figure 7(f)). If we have only a line of 3 meta-modules, it suffices to contract and expand one meta-module, respectively (see the second round in Figure 7(f)). Note that we have to remove most of the bonds along the line to permit the line to move freely. The pull and push operations allow us to transfer $R$ from one corner to the next corner in a single round.
@@
In particular, note that a $1$-tunnel allows us to move a meta-module around another one (see Figure 7(d)). In other models for modular robot systems, this simple case is known as the rotation primitive.
Lemma 3.6.
Our implementation of the rotation primitive requires $2$ rounds.
Akitaya et al. [2] have introduced two further movement primitives for rectangular MRSs: the leapfrog and the monkey (see Figure 9). In these primitives, the moving module performs a ‘jump’ between two not adjacent modules. We are able to transfer these primitives to our rhombical meta-modules. The idea is to divide the meta-modules into submodules to bridge the gap between the start and target position of the moving meta-module (see Figure 10).
Lemma 3.7.
@@
Figure 9: Leapfrog and monkey jumps. These movement primitives were introduced by Akitaya et al. [2]. We can simulate these by dividing the meta-modules into submodules. With these, we build a bridge to the new position of the meta-module by applying slide and $k$-tunnel primitives.
(a) Leapfrog jump.
(b) Straight monkey jump.
(c) Diagonal monkey jumps.
Figure 10: Leapfrog and monkey jumps. These movement primitives were introduced by Akitaya et al. [2]. We can simulate these by dividing the meta-modules into submodules. With these, we build a bridge to the new position of the meta-module by applying slide and $k$-tunnel primitives.
@@ 3.2 Hexagonal Meta-Modules
Let $\ell$ be an even integer as before. Our hexagonal meta-module consists of three rhombical meta-modules of side length $\ell-1$ (see Figure 10(a)). We arrange them into a hexagon of alternating side lengths $\ell$ and $\ell-1$.
There are two possibilities to arrange the rhombical meta-modules in the hexagon (see Figure 10(a)). We can switch between them as follows. The hexagonal meta-module can be split into three equilateral triangles of side length $\ell-1$, and three equilateral triangles of side length $\ell-2$. The idea is to apply a similar technique as in the realignment primitive for rhombical meta-modules. In the first round, each second row contracts but without shifting the other rows. In the second round, the rhombi interchange the smaller triangles through handovers. In the third round, we expand the contracted amoebots into the opposite direction – again without shifting the other rows.
@@
Our implementation of the switching primitive requires $3$ rounds.
(a) Switching primitive.
(b) Rotation primitive.
Figure 11: Movement primitives for hexagonal meta-modules.
We can arrange the meta-modules on a hexagonal tesselation of the plane (see Figure 0(c)). In the following, we introduce the rotation primitive for hexagonal meta-modules.
In the rotation primitive, we move a hexagonal meta-module $H_{1}$ around another hexagonal meta-module $H_{2}$ as follows (see Figure 10(b)). We arrange $H_{2}$ such that a rhombical meta-module $R_{2}$ is adjacent to both the old and new position of $H_{1}$, and $H_{1}$ such that the rhombical meta-module $R_{1}$ adjacent to $R_{2}$ is aligned to the same axes as $R_{2}$. We contract $R_{1}$ and $R_{2}$, and then expand them into the direction of the new position of $H_{1}$ (compare to Section 3.1). This movement primitive requires two rounds. Note that additional steps may be necessary to switch or reorientate the rhombical meta-modules beforehand.

### v3

@@ 2.3 Problem Statement and Our Contribution
In this paper, we formalize and extend the joint movement extension proposed by Feldmann et al. [23]. In the following, we provide a proof of concept. For that, we focus on two fundamental problems of MRSs: reconfiguration and locomotion. We study these problems from a centralized view to explore the limits of the extension.
In the reconfiguration problem, an MRS has to reconfigure its initial structure into a given target structure. For that, we define meta-modules of rhombical and hexagonal shape. We show that these meta-modules are able to perform various movement primitives of other MRSs, e.g., crystalline atoms, and rectangular/hexagonal metamorphic robots. This allows us to simulate the reconfiguration algorithms of those models.
@@ 2.4 Related Work
Various MRSs have been defined for different lattices, e.g., [12, 62] utilize the triangular grid, and [3, 22, 51] utilize the Cartesian grid. We will build meta-modules for the Cartesian and triangular grid. Note that some MRSs were also physically realized. Examples for MRSs using the triangular grid are hexagonal metamorphic robots [12], HexBots [52], fractal machines [43], and catoms [35]. Examples for MRSs using the Cartesian grid are CHOBIE II [56], EM-Cubes [7], M-Blocks [49], pneumetic cellular robots [29], and XBots [61].
@@
The most common movement primitives for MRSs on the Cartesian grid are the rotation and slide primitives (see Figures 8(b) and 8(c)), and for MRSs on the triangular grid the rotation primitive (see Figure 8(a)). Some models may constrain these movement primitives (see Figure 9). We refer to [2, 28] for a deeper discussion about possible constraints. One way to bypass such constraints is to construct meta-modules, e.g., see [19, 28, 44]. Our meta-modules implement all aforementioned movement primitives without any constraints, i.e., they perform them in place.
@@
(f) Leapfrog jump.
(g) Straight monkey jump.
(h) Diagonal monkey jumps.
@@
Figure 9: Constrained movement primitives. The green modules are moving. The gray cells must be empty. Both primitives have the same result while the pivoting model requires more free space than the sliding model.
From all aforementioned models, crystalline atoms [50], telecubes [55, 59], and prismatic cubes [60] are the closest to our joint movement extension. In these MRSs, each robot has the shape of a unit square and is able to expand an arm from each side by half a unit. Adjacent robots are able to attach and detach their arms. Similar to our joint movement extension, a robot can move attached robots by expanding or contracting its arms. In contrast to amoebots, a line of robots cannot reconfigure to any other shape since each pair of robots can only have a single point of contact [5]. In order to allow arbitrary reconfigurations, the robots are combined into meta-modules of square shape that are capable of various movement primitives, e.g., the rotation, slide, contraction, and tunnel primitives (see Figures 8(b), 8(d), 8(c) and 8(e)). Our rhombical meta-modules can implement all these movement primitives. The rhombical shape has two advantages compared to the square shape. First, we can implement further movement primitives that provide us with a simple way to construct a walking structure. Second, we can utilize rhombical meta-modules as a basis for hexagonal meta-modules.
@@ 3 Meta-Modules
In this section, we will combine multiple amoebots to meta-modules. In other models for programmable matter and modular robots, meta-modules have proven to be very useful. For example, they allow us to bypass restrictions on the reconfigurability [19, 58] and to simulate (reconfiguration) algorithms for other models [4, 48].
In the first two subsections, we will present meta-modules of rhombical and hexagonal shape, respectively. In the third subsection, we show how to construct such meta-modules from an arbitrary amoebot structure. In the fourth subsection, we discuss how to increase the stability of the structure during the movements.
@@ 3.1 Rhombical Meta-Modules
Let $\ell$ be a positive even integer. Our rhombical meta-module consists of $\ell^{2}/2$ uniformly oriented expanded amoebots that we arrange into a rhombus of side length $\ell-1$ (see the left side of Figure 10). We obtain a parallelogram of side lengths $\ell-1$ and $\ell/2-1$ if we contract all amoebots (see Figure 10). Note that we have to remove some bonds to perform the contraction. We can expand the parallelogram again by reversing the contractions.
Figure 10: Contraction primitive.
@@
There are exactly two possibilities to arrange the uniformly oriented expanded amoebots in a rhombus. We can reorientate all amoebots within a rhombus as follows (see Figure 11). First, we partition the meta-module into $\frac{\ell}{2}\times\frac{\ell}{2}$ submodules. Note that each submodule consists of exactly $2$ amoebots. Then, we color the submodules in a checkerboard pattern. Finally, we reorientate both colors independently of each other within two phases. Each phase consists of two rounds. In the first round, we contract each submodule of one color into its shorter diagonal. The submodules of the other color ensure that the meta-module remains connected. In the second round, we expand all contracted amoebots into the other corner of its submodule.
Figure 11: Reorientation primitive.
@@
Our implementation of the reorientation primitive requires $4$ rounds.
Furthermore, there are three possibilities to align the sides of a rhombus to the axes of the triangular grid. By sliding each second row along its axis to the other side, we can realign the other axis of the rhombus to the third axis of the triangular grid (see Figure 12). Note that in combination with the reorientation primitive, we are able to align a rhombus with any two axes of the triangular grid.
Figure 12: Realignment primitive.
@@
We can arrange the meta-modules on a rhombical tesselation of the plane if they are all aligned to the same axes (see Figure 13). Note that due to the triangular grid, the meta-modules are not connected diagonally everywhere. Hence, we will only consider meta-modules connected if their sides are connected.
Figure 13: Rhombical tesselation. Note that there are spaces between the meta-modules since a node cannot be occupied by more than one amoebot. However, the spaces do not contain any nodes.
In the following, we introduce three movement primitives: the slide, rotation, and $k$-tunnel primitive. Our implementations of these primitives are similar to the ones for crystalline robots (e.g., [5]) and teletubes (e.g., [59]).
In the slide primitive, we move a meta-module $R_{1}$ along two adjacent substrate meta-modules $R_{2}$ and $R_{3}$ (see Figure 8(c)). We realize the primitive as follows (see Figure 14). We assume that all amoebots are orientated into the movement direction. Otherwise, we apply the reorientation primitive. With respect to Figure 14, let $L_{1}$ denote the uppermost layer of $R_{2}$ and $R_{3}$, and $L_{2}$ the second uppermost layer of $R_{2}$ and $R_{3}$. Our slide primitive consists of two rounds. In the first round, we contract all amoebots in $L_{1}$ after removing all bonds between $L_{1}$ and $R_{1}$ except the last one in the movement direction, and all bonds between $L_{1}$ and $L_{2}$ except the last one in the opposite direction. This moves $R_{1}$ into its target position. In the second round, we restore $R_{2}$ and $R_{3}$. For that, we expand $L_{1}$ again after removing all bonds between $L_{1}$ and $R_{1}$, and between $L_{1}$ and $L_{2}$ except the last ones in the movement direction, respectively. This ensures that $R_{1}$ stays in place.
Figure 14: Slide primitive.
@@
In the rotation primitive, we move a meta-module $R_{1}$ around another meta-module $R_{2}$ (see Figure 8(b)). We realize the primitive as follows (see Figure 15). We assume that all amoebots are orientated towards each other. Otherwise, we apply the reorientation primitive. Our rotation primitive consists of two rounds. In the first round, we pull $R_{1}$ into $R_{2}$ by contracting both meta-modules into parallelograms. The resulting contracted meta-modules form a rhombus consisting of $\ell^{2}$ contracted amoebots. Since all amoebots are contracted, we can rotate $R_{1}$ and $R_{2}$ within that rhombus by exchanging amoebots. This allows us in the second round to push $R_{1}$ out of $R_{2}$ into $R_{1}$’s target position.
Figure 15: Rotation primitive for rhombical meta-modules. Red meta-modules perform a pull operation, and blue meta-modules a push operation.
Lemma 3.5.
Our implementation of the rotation primitive requires $2$ rounds.
In the $k$-tunnel primitive, we move a meta-module $R$ through a simple path of meta-modules with $k$ corners to the other end (see Figure 8(e)). We realize the primitive as follows (see Figure 16). We assume that all meta-modules in the corners are orientated towards the preceding meta-module, and all other meta-modules are oriented along the path. Otherwise, we apply the reorientation primitive. We will not move $R$ directly through the path. Instead, we will make use of pull and push operations (see Figure 15).
Consider a line of at least 4 meta-modules with two contracted meta-modules at one end. By expanding those two meta-modules and contracting the two meta-modules at the other end, we transfer a meta-module from one end to the other end without changing the length of the line (see the third round in Figure 16). If we have only a line of 3 meta-modules, it suffices to contract and expand one meta-module, respectively (see the second round in Figure 16). Note that we have to remove most of the bonds along the line to permit the line to move freely. The pull and push operations allow us to ‘transfer’ $R$ from one corner to the next corner in a single round. Note that the rotation primitive is a special case of the $k$-tunnel primitive.
Figure 16: $k$-tunnel primitive. Red meta-modules perform a pull operation, and blue meta-modules a push operation.
@@
Akitaya et al. [2] have introduced two further movement primitives for rectangular MRSs: the leapfrog and the monkey primitive (see Figures 8(f), 8(g) and 8(h)). In these primitives, the moving module performs a ‘jump’ between two not adjacent modules. We are able to transfer these primitives to our rhombical meta-modules. The idea is to divide the meta-modules into submodules and then build a bridge between the start and target position of the moving meta-module by applying slide and $k$-tunnel primitives (see Figure 17).
@@
Figure 17: Leapfrog and monkey jump primitives. Each step consists of multiple rounds.
Lemma 3.7.
@@ 3.2 Hexagonal Meta-Modules
Let $\ell$ be an even integer as before. Our hexagonal meta-module consists of three rhombical meta-modules of side length $\ell-1$ (see Figure 18). We arrange them into a hexagon of alternating side lengths $\ell$ and $\ell-1$.
There are two possibilities to arrange the rhombical meta-modules in a hexagonal meta-module (see Figure 18). We can switch between them as follows. We assume that all amoebots are orientated as shown in Figure 18. Otherwise, we apply the reorientation primitive. Our implementation requires two rounds. However, for the sake of comprehensibility, we will split both rounds into two rounds, respectively.
The idea is to apply a similar technique as in the realignment primitive for rhombical meta-modules. In the first round, each rhombical meta-module contracts each second row without shifting the other rows. Each rhombical meta-module within the hexagonal meta-module can be split into an equilateral triangle of side length $\ell-1$, and an equilateral triangle of side length $\ell-2$. In the second and third round, the rhombical meta-modules exchange the smaller triangles. For that, we first contract all expanded amoebots in each smaller triangle such that the amoebots form lines between the adjacent bigger triangles. Note that the amoebots in the corners of the hexagonal meta-module ensure connectivity to adjacent hexagonal meta-modules. Then, we expand the other amoebots in each smaller triangle into the nodes previously occupied by the expanded amoebots. In the fourth round, each resulting rhombical meta-module expands each second row into the opposite direction (of the contracted amoebots of its bigger triangle) – again without shifting the other rows. Finally, note that we can combine the first and second round resp. the third and fourth round. For that, we have to remove all necessary bonds at once, respectively.
Figure 18: Switching primitive.
@@
Our implementation of the switching primitive requires $2$ rounds.
We can arrange the meta-modules on a hexagonal tesselation of the plane (see Figure 19). In the following, we introduce the rotation primitive for hexagonal meta-modules.
Figure 19: Hexagonal tesselation. Note that there are spaces between the meta-modules since a node cannot be occupied by more than one amoebot. However, the spaces do not contain any nodes.
In the rotation primitive, we move a hexagonal meta-module $H_{1}$ around another hexagonal meta-module $H_{2}$ as follows (see Figures 8(a) and 20). We arrange $H_{2}$ such that a rhombical meta-module $R_{2}$ is adjacent to both the old and new position of $H_{1}$, and $H_{1}$ such that the rhombical meta-module $R_{1}$ adjacent to $R_{2}$ is aligned to the same axes as $R_{2}$. We contract $R_{1}$ and $R_{2}$, and then expand them into the direction of the new position of $H_{1}$ (compare to Lemma 3.5). This movement primitive requires two rounds. Note that additional steps may be necessary to switch or reorientate the rhombical meta-modules beforehand.
Figure 20: Rotation primitive for hexagonal meta-modules.
@@
3.3 Meta-Module Construction
First, we form a line of $r$ rhombical meta-modules of side length $\ell-1$ from a line of $n=r\cdot\ell^{2}/2$ amoebots as follows (see Figure 21). First, we divide the line of amoebots into groups of $\ell$ amoebots. The idea is to apply the realignment primitive (see Lemma 3.3) to rotate each second group once counterclockwise by $60^{\circ}$ and each other group twice clockwise by $60^{\circ}$ without breaking the connectivity along the line. However, in order to apply the realignment primitive, we first have to expand the groups into parallelograms of side lengths $\ell-1$ and $1$.
Therefore, the procedure consists of the following seven phases. In the first six phases, we alternate between two types of phases. In the first type, we possibly remove the parallelograms of the previous phase (third, and fifth phase) and form the parallelograms for the next phase (first, third, and fifth phase). In the second type, we apply the realignment primitive to rotate the parallelograms by $60^{\circ}$ (second, fourth, and sixth phase). Finally, in the seventh phase, we obtain the rhombical meta-modules by expanding all contracted amoebots. We obtain the following theorem.
Figure 21: Construction of rhombical meta-modules. Adjacent amoebots of the same color belong to the same group. Each row shows a phase.
Theorem 3.10.
There is a centralized algorithm that forms a line of $r$ rhombical meta-modules of side length $\ell-1$ from a line of $n=r\cdot\ell^{2}/2$ amoebots within $12$ rounds.
Proof 3.11.
Each removal and formation requires one round. Hence, the first phase requires a single round, and the third and fifth phase require two rounds, respectively. By Lemma 3.3, each phase of the second type requires two rounds. The seventh phase requires a single round since we only expand all contracted amoebots. Therefore, the algorithm requires $12$ rounds overall.
Remark 3.12.
Theorem 3.10 also improves a result by Feldmann et al. [23]. With the help of handovers, they showed how to reconfigure a line of amoebots into a rhombus within $O(\log n)$ rounds. For $r=1$, we obtain the same result within $12$ rounds without using handovers.
Second, we form a line of $h=\frac{r}{3}$ hexagonal meta-modules of alternating side lengths $\ell$ and $\ell-1$ from a line of $r$ rhombical meta-modules of side length $\ell-1$ as follows (see Figure 22). We assume that all rhombical meta-modules are oriented along the line. Otherwise, we apply the reorientation primitive. We first divide the line of rhombical meta-modules into groups of $3$ rhombical meta-modules and apply the following three steps to each group.
In the first step, we apply the rotation primitive to rotate the first meta-module around the second one. In the second step, we apply the reorientation primitive to the first and third meta-module such that in the third step, we can apply the realignment primitive to the them to form a hexagonal meta-module. We obtain the following theorem.
Figure 22: Construction of hexagonal meta-modules. Adjacent meta-modules of the same color belong to the same group. The first two rounds show the application of the rotation primitive. The last two rounds show the application of the realignment primitive. We have omitted the application of the reorientation primitive.
Theorem 3.13.
There is a centralized algorithm that forms a line of $h=\frac{r}{3}$ hexagonal meta-modules of alternating side lengths $\ell$ and $\ell-1$ from a line of $r$ rhombical meta-modules of side length $\ell-1$ within $8$ rounds.
Proof 3.14.
By Lemma 3.5, the first step requires $2$ rounds. By Lemma 3.2, the second step requires $4$ rounds. By Lemma 3.3, the third step requires $2$ rounds. Therefore, the algorithm requires $8$ rounds overall.
Corollary 3.15.
There is a centralized algorithm that forms a line of $h$ hexagonal meta-modules of alternating side lengths $\ell$ and $\ell-1$ from a line of $n=3\cdot h\cdot\ell^{2}/2$ amoebots within $20$ rounds.
Proof 3.16.
The corollary directly follows from Theorems 4.3 and 4.5. Note that all rhombical meta-modules are oriented along the line after applying Theorem 4.3 such that no additional reorientations are necessary.